OpenStudy (anonymous):

Find the value by referring to the graph of the sine function. sin(-7/2 pi)

OpenStudy (anonymous):

Do you not know how to do the math, or do you simply not understand the task? It's horribly simple with a calculator, just type in the equation. But of course that's not what you have to do. Also, is it \(sin(\frac{7}{2}\pi)\), or \(sin(\frac{7}{2\pi})\) ?

OpenStudy (anonymous):

I can't just put it into the calculator to solve it or ill get a completely different answer than what I need. I've tried it like that. It has something to do with the graph and idk how to do it. it's \[\sin (-\frac{ 7 }{ 2 }\pi)\]